### v1

# Cherry on Top: Parameter Heterogeneity and Quantization in Large Language Models

Abstract
This paper reveals the phenomenon of parameter heterogeneity in large language models (LLMs). We find that a small subset of ”cherry” parameters exhibit a disproportionately large influence on model performance, while the vast majority of parameters have minimal impact. This heterogeneity is found to be prevalent across different model families, scales, and types. Motivated by this observation, we propose CherryQ, a novel quantization method that unifies the optimization of mixed-precision parameters. CherryQ identifies and preserves the critical cherry parameters in high precision while aggressively quantizing the remaining parameters to low precision. Extensive experiments demonstrate the effectiveness of CherryQ. CherryQ outperforms existing quantization approaches in terms of perplexity and downstream task performance. Notably, our 3-bit quantized Vicuna-1.5 exhibits competitive performance compared to their 16-bit counterparts. These findings highlight the potential of CherryQ for enabling efficient deployment of LLMs by taking advantage of parameter heterogeneity.
1 Introduction
The rapid development of large language models (LLMs) has increased the demand of efficient deployment in various environments [1, 23, 12, 2]. However, the parameter size poses significant challenges for GPU memory requirements. Quantization, which reduces the bit-width of model parameters, has emerged as a solution to alleviate memory constraints of LLM deployment [14, 13, 24, 25, 18].
Quantizing parameters from higher bits to the integer points in the lower-bit space inevitably perturbs the parameters from its optimum, leading to a degradation in performance (i.e. quantization error). To mitigate this error, various approaches have been proposed, such as iterative block-wise quantization ([15, 10, 8]), gradient approximation ([18, 3]), and low-rank approximation of perturbations ([9]). However, existing approaches still cause clear performance degradation, especially for extreme low bits (e.g. 3-bit).
Before investigating how to further mitigate the performance degradation, we raise a more fundamental question: To what extent can quantization errors be mitigated? Our study shows that the answer is more complex than expected. For the vast majority ($>99\%$) of parameters, their quantization errors are minimal and thus can be alleviated or ignored. Nonetheless, there exists a small subset of parameters ($<1\%$) for which the quantization errors are substantial and hard to mitigate.
(a) LLaMA2 7B
layers.15.self_attn.v_proj
(b) LLaMA2 13B
layers.15.self_attn.v_proj
(c) Mistral 7B
layers.15.self_attn.q_proj
(d) Gemma 7B
layers.25.mlp.down_proj
(e) Vicuna-1.5 7B
layers.15.self_attn.v_proj
(f) Vicuna-1.5 13B
layers.16.self_attn.v_proj
Figure 1: Scatter plot of parameter impacts in different LLMs. We randomly sampled 4096 parameters from the corresponding parameter matrix. Each point represents the impact of an individual parameter. Insets show the zoomed-in y-axis. The heterogeneity is found in different model scales (1(a),1(b)), different model families (1(c), 1(d)), and both base models and chat models (1(e), 1(f)).
Consider Figure 1(a) as an example. We show the scatter plot of impacts on quantization error when perturbing each individual parameters in a parameter matrix from LLaMA2-7b [23]. The derivation of impacts are detailed in § 3. As 99% of parameters are in the range of (0,0.1), a small subset of “cherry” parameters exhibit a disproportionately large influence in the range of (5,30), with 50-300 times greater than the maximum value of the remaining 99% parameters.
This phenomenon is not an isolated occurrence. We observed similar patterns across different LLM scales (Figure 1(a)1(b)), different LLM families, including Mistral [12] (Figure 1(c)) and Gemma [22] (Figure 1(d)), and both base models and chat models (Vicuna-1.5 [5] Figure 1(e)1(f)). The consistent presence suggests that it is an inherent characteristic of LLMs. We will elaborate the evidences from a more macro view in § 5. Based on these findings, we introduce the phenomenon of parameter heterogeneity:
Parameter Heterogeneity in LLMs:
Consider the impact of parameter perturbations on the model’s behavior, a small set of “cherry” parameters have a disproportionately large influence. In contrast, the vast majority of normal parameters have a minimal impact.
We emphasize that the parameter impact is heterogeneous, as it exhibits a more severe imbalance compared to the commonly observed imbalance in LLMs [18, 17, 6, 7]. As a contrast, in § 5, we demonstrate that the parameter impact is much more heterogeneous than the parameter magnitude, although the latter one is also found to be imbalanced [7, 18]. This stark difference highlights the significance of considering parameter impact when optimizing and quantizing LLMs.
The parameter heterogeneity poses new challenges for conventional quantization strategies. Although quantizing most normal parameters will cause little performance degradation, cherry parameters are highly sensitive to quantization-induced perturbations. However, existing strategies usually fail to preserve the delicate structure of these critical parameters [8, 18].
Addressing this challenge is not trivial. One straightforward way is to employ a mixed-precision quantization [13], representing cherry parameters and normal parameters with high and low precisions, respectively. However, the simultaneous optimization of both types of parameters becomes a major challenge. In the widely used GPTQ [8] approach and PTQ framework, the optimal values of the early quantized parameters may change as the cherry parameters are updated. However, parameters cannot be updated once they are quantized in PTQ, which limits the early quantized parameters from reaching their optimal values.
To address this challenge, we employ the quantization-aware training (QAT) framework to handle parameter heterogeneity. Our quantization optimize the mixed-precision parameters with a unified backpropagation. For the cherry parameters, we maintain their 16-bit representations and apply standard gradient descent. For the normal parameters, we apply an extra Straight-Through Estimator (STE) trick [3] for gradient descent. Therefore, all parameters are continuously and uniformly optimized. We denote the approach as CherryQ (Cherry parameter and Quantization-aware training).
Extensive experiments on different models and benchmarks demonstrate the efficacy of CherryQ. It consistently yields the lowest perplexity on most settings. Notably, our 3-bit Vicuna-1.5 model exhibit performance on par with the 16-bit counterpart on Vicuna-bench [5].
We believe that our work opens up new avenues for understanding and harnessing the complex interplay of parameters in LLMs. The heterogeneity offers a novel perspective to navigate the trade-off between parameter efficiency and performance.
2 Related Work
Quantization Strategies for LLMs Various quantization strategies have been proposed in the literature to reduce the precision of weights and activations while maintaining acceptable accuracy. These strategies can be broadly categorized into post-training quantization and quantization-aware training [14]. Post-training quantization methods, such as OBD, OBS, and GPTQ, directly quantize the pre-trained model without fine-tuning [15, 10, 8]. On the other hand, quantization-aware training methods, such as LLM-QAT [18], incorporate quantization operations into the training process to jointly optimize the quantized model. Some works also explore mixed-precision quantization [13] and adaptive quantization bins [7] to achieve a better trade-off between accuracy and efficiency.
Outliers in Language Model Quantization The idea of modeling parameter outliers in LLM quantization is not new. The exploration of outliers primarily includes the perspectives of magnitude [18, 7] and activations [4, 6]. For example, from the magnitude perspective, QLoRA assumes that parameters follow a Gaussian distribution [7] and designs information-theoretically optimal quantized bins based on this assumption. [18] keeps outlier parameters in 16-bit precision. From the activation perspective, [17] migrates the outlier amplifier to subsequent modules through an equivalent transformation. Additionally, SqueezeLLM also measures outliers from the perspective of parameter impact [13]. To the best of our knowledge, our work is the first to systematically reveal the outliers (heterogeneity) of parameter impact across different models, and we show a more pronounced imbalance in parameter impacts compared to magnitudes (§ 6.4). Furthermore, we propose a method to unify outlier (cherry) parameter optimization and normal parameter optimization, addressing the optimization challenges of heterogeneous parameters.
3 Quantifying the Impact of Parameters on Model Performance
The impact of parameters on model performance is quantified by the increase of the training loss when perturbing the parameter weight, which is widely used in post-training quantization approaches [15, 10, 8]. We adopt a second-order Taylor approximation of the training loss w.r.t. parameter perturbation. Given a parameter $w_{i}$ and a small perturbation $\Delta$ applied to it, such that $w_{i}\leftarrow w_{i}+\Delta$, the change in the training loss can be expressed as:
$\mathbf{L}(w_{i}+\Delta)-\mathbf{L}(w_{i})=g_{i}\Delta+\frac{1}{2}\mathbf{H}_{ii}\Delta^{2}+O(|\Delta|^{3})$ (1)
where $g_{i}=\mathbb{E}[\frac{\partial L}{\partial w_{i}}]$ represents the expected gradient of the loss with respect to $w_{i}$, and $\mathbf{H}_{ii}=\mathbb{E}[\frac{\partial^{2}L}{\partial w_{i}^{2}}]$ denotes the $i$-th value of the Hessian matrix of the loss. Since the target model is a well-converged model, we can assume that $g_{i}\approx 0$, simplifying the expression to:
$\mathbf{L}(w_{i}+\Delta)-\mathbf{L}(w_{i})\approx\frac{1}{2}\mathbf{H}_{ii}\Delta^{2}$ (2)
Therefore, $\mathbf{H}_{ii}$ quantify the impact of quantization-induced perturbations on the model’s training loss. Parameters with larger values of $\mathbf{H}_{ii}$ exhibit higher sensitivity to quantization and require careful treatment to maintain model performance. We denote $\mathbf{H}_{ii}$ as the impact of $w_{i}$.
Efficient Computation Computing $\mathbf{H}_{ii}$ of the diagonal of Hessian matrix for each parameter is computationally expensive, particularly for large-scale models. To overcome this challenge, we propose an efficient approximation using the Fisher Information Matrix ($\mathbf{F}$). Since $\mathbf{H}$ is the Hessian matrix of a negative log-likelihood loss, $\mathbf{H}$ is equal to Fisher information matrix [16]. For the diagonal of the Hessian matrix, we have:
$\mathbf{H}_{ii}=\mathbf{F}_{ii}=\mathbb{E}[g_{i}^{2}]$ (3)
4 Unified Mixed-Precision Training
The insights gained from Figure 1 highlights the heterogeneity in model parameters. The cherry parameters, despite constituting less than 1% of the total parameter count, exert a substantial influence on the model. Indiscriminately quantizing these cherry parameters alongside the normal parameters may lead to a significant deterioration in model performance.
To mitigate the impact of cherry parameters on quantization, we propose to preserve their high-precision values during the quantization process. By maintaining the fidelity of these critical parameters, we ensure that the essential information they capture is not compromised.
Optimizing mixed-precision parameters in LLMs presents a unique challenge. The widely adopted GPTQ approach [8], which falls under the Post-Training Quantization (PTQ) framework [14], struggles to simultaneously optimize high-precision cherry parameters and low-precision normal parameters. This is because updating the cherry parameters during the PTQ process significantly affects the model, causing the optimal values of the normal parameters to vary. However, in the PTQ framework, once the parameters are quantized, they cannot be updated further. This limitation prevents the early-stage quantized parameters from reaching their optimal values. On the other hand, if we do not allow the updates of the cherry parameters during the PTQ process [17], the quantized model will lose the flexibility provided by these critical parameters.
To address this challenge, we propose a novel approach that unifies the optimization of mixed-precision parameters. Our method leverages a QAT framework, which allows for the simultaneous optimization of both cherry parameters and normal parameters. During backpropagation, the high-precision cherry parameters are updated using standard gradient descent, while the low-precision normal parameters employ the Straight-Through Estimator (STE) trick [3] for low precision gradient descent. This unified backpropagation enables end-to-end optimization of both cherry parameters and normal parameters, enhancing the overall optimization effect. We show the quantization in Algorithm 1.
Algorithm 1 Cherry Parameter and Quantization-aware Training (CherryQ)
1:Model parameters $\mathbf{W}$, quantization function $Quant(\cdot)$, threshold $\tau$, learning rate $\eta$
2:Quantized model parameters
3:$\mathbf{C}\leftarrow\{w_{i}\in\mathbf{W}\mid H_{ii}>\tau\}$triangleright Identify cherry parameters
4:$\mathbf{N}\leftarrow\mathbf{W}\setminus\mathbf{C}$triangleright Identify normal parameters
5:for each training batch $x$ do
6:     $L\leftarrow\mathrm{model}(x;\mathbf{C}\cup Quant(\mathbf{N}))$triangleright Compute loss
7:     $\mathbf{C}\leftarrow\mathbf{C}-\eta\frac{\partial L}{\partial\mathbf{C}}$triangleright Standard gradient descent
8:     $\mathbf{N}\leftarrow\mathbf{N}-\eta\cdot\mathrm{STE}(\frac{\partial L}{\partial\mathbf{N}})$triangleright Gradient descent with gradient approximation by STE
9:end for
10:return $\mathbf{C}\cup Quant(\mathbf{N})$
5 Prevalence of Parameter Heterogeneity in LLMs
While Figure 1 showcases the heterogeneity of selected parameter matrices in different LLMs, it is crucial to investigate whether this phenomenon is pervasive across the hundreds of parameter matrices within each LLM. In this section, we conduct a comprehensive analysis of parameter heterogeneity from a macro perspective.
To quantify the degree of heterogeneity in a parameter matrix, we introduce the heterogeneity score of the matrix. Inspired by the observation in Figure 1, where a small subset of parameters exhibits significantly higher impacts compared to the maximum of the majority, we define the heterogeneity score as the ratio of the mean impact of the top 1% parameters to the maximum impact of the bottom 99% parameters, as shown in Equation (4). A higher heterogeneity score indicates a more pronounced disparity in parameter importance within the matrix.
Heterogeneity Score=Mean($\text{Heterogeneity Score}=\frac{\text{Mean}(\mathbf{H}_{ii\text{ top 1\%}})}{\text{Max}(\mathbf{H}_{ii\text{ bottom 99\%}})}$ (4)
For comparison, we also include the heterogeneity scores based on the magnitude of parameters, a commonly used measure of parameter importance [11]. The magnitude-based heterogeneity score is calculated using Equation (5).
Magnitude Heterogeneity $\text{Magnitude Heterogeneity Score}=\frac{\text{Mean}(|w_{i}|_{\text{ top 1\%}})}{\text{Max}(|w_{i}|_{\text{ bottom 99\%}})}$ (5)
To provide a comprehensive view of parameter heterogeneity across different matrices, we plot the scatter distribution of heterogeneity scores for all parameter matrices of each model in Figure 2. It clearly reveals that the parameter matrices across different LLMs exhibit high heterogeneity scores, especially when comparing with parameter magnitudes. This finding strongly suggests that parameter heterogeneity is not an isolated occurrence but rather a widespread phenomenon in LLMs.
The pervasiveness of parameter heterogeneity highlights the need for quantization strategies that can effectively handle the disparate importance of parameters, ensuring that the cherry parameters are preserved with higher precision while allowing for more aggressive quantization of the less influential normal parameters.
(a) LLaMA-2 7B
(b) LLaMA-2 13B
(c) Mistral 7B
(d) Gemma 7B
(e) Vicuna-1.5 7B
(f) Vicuna-1.5 13B
Figure 2: Scatter distribution of heterogeneity scores for different parameter matrices in LLMs. Each point represents a parameter matrix, with the x-axis indicates the matrix index and the y-axis showing the heterogeneity score.
6 Quantization Experiments
In the experimental section, we demonstrate the effectiveness of the CherryQ for both base LLMs and chat LLMs. We also compare different cherry parameter selection criteria to highlight the impact-based heterogeneity.
6.1 Implementation Details
Parameter Representation: Based on the observation that cherry parameters occupy a very small proportion, for each row of parameters in each parameter matrix, we only consider the top 1/256 parameters with the highest impact as cherry parameters and retain their FP16 precisions. For example, the parameter matrix size of LLaMA2-7B is $4096\times 4096$. So we select $16$ parameters with the highest impact for each row, which results in $4096\times 16$ parameters as cherry parameters. Additionally, to recover the complete parameter matrix, an INT16 is required to record the column index for each cherry parameter. So each cherry parameter need 32 bits.
For normal parameters, we employ full range symmetric MinMax quantization to quantize their weights [14]. Specifically, an FP16 value is mapped to the range of $[-2^{k-1},2^{k-1}-1]$ and symmetrically distributed on both sides of the coordinate axis. The quantization of an FP16 tensor $X^{FP16}$ to $k$ bits is computed by:
$X^{Intk}=\lfloor Clip(\frac{X^{FP16}}{S},-2^{k-1}+\epsilon,2^{k-1}-\epsilon)-0.5\rceil$ (6)
[…]
Dequantization restores the quantized integer values based on the scaling factor:
$Dequant(S,X^{Intk})=S(X^{Intk}+0.5)$ (7)
To further improve the quantization accuracy, we adopt a widely-used parameter grouping strategy. Specifically, the parameters are divided into groups in order, and each group independently calculates its scaling factor. For example, if we divide a parameter matrix $W\in\mathbb{R}^{r\times c}$ that needs to be quantized with a group size of $B$, we will obtain a total of $r\times(c/B)$ groups.
Quantization Datasets: For the quantization of the base LLMs, we follow [8] to use C4 [20] as the training data. We select the first 4 partitions of C4 and choose data with a length of $\geq 2048$ tokens, resulting in a total of 50k samples of 2048 token. For the chat LLMs, since Vicuna-1.5 [5] is obtained by supervised fine-tuning based on ShareGPT [5], we also use the ShareGPT dataset for training. We utilize a total of 20k training samples from ShareGPT for QAT and Cherry.
Baselines We compare our method with various quantization methods, including QAT [18], GPTQ [8], SqueezeLLM [13], OminiQuant [21], and AWQ [17]. For OminiQuant and AWQ, we use their results reported in [21]. For SqueezeLLM, we use the results in its original paper [13]. For GPTQ, its 4-bit model is obtained from the open-source  [^1] https://huggingface.co/TheBloke. Due to the lack of a 3-bit GPTQ model, we quantize the model ourselves via the implementation of Auto-GPTQ  [^2] https://github.com/AutoGPTQ/AutoGPTQ. Since CherryQ is based on QAT, for fair comparisons, the implementation of QAT is the same as CherryQ, except that it does not handle cherry parameters.
6.2 Effect of Base LLM Quantization
In this section, we present the main experimental results demonstrating the effectiveness of CherryQ on LLaMA2 [23]. We evaluate CherryQ with both perplexity and downstream tasks, comparing its performance against state-of-the-art quantization methods.
6.2.1 Perplexity Results
We follow [8, 21] to evaluate the perplexity of CherryQ on two widely-used corpora: C4 and WikiText-2 [19]. We use the validation split of C4 to avoid data leakage. We show the results of 3-bit quantization using different quantization approaches in Table 1. We show the results of different model scales and different group sizes.
From the results, CherryQ consistently outperforms all other approaches across both model sizes (7B and 13B) and grouping sizes (64 and 128), achieving the lowest perplexity on both the C4 and WikiText-2 datasets. Notably, CherryQ’s perplexity is significantly closer to the full-precision (FP16) baseline compared to other methods, highlighting its ability to preserve model performance after quantization.
Table 2 compares different 4-bit quantization methods. Again, CherryQ achieves the lowest perplexity scores in most settings, demonstrating its effectiveness in higher-bit quantization settings.
6.2.2 Downstream Task Performance
To further validate the effectiveness on specific tasks, we evaluate the quantized models on various downstream tasks from the HuggingFace OpenLLM Leaderboard. Table 3 presents the performance comparison of different 3-bit quantization methods for LLaMA2. CherryQ consistently outperforms other methods across almost all tasks, achieving the highest average score. This showcases CherryQ’s ability to maintain the model’s generalization capabilities for downstream tasks.
Table 1: Perplexity ($\downarrow$) of 3-bit quantization on LLaMA2 models . gX means the group size is X. The results of OminiQuant and AWQ are from [21]. The results of SqueezeLLM are from [13].
Method 7B-3bit-g128 7B-3bit-g64 13B-3bit-g128 13B-3bit-g64
c4 wiki2 c4 wiki2 c4 wiki2 c4 wiki2
FP16 6.97 5.47 6.97 5.47 6.47 4.88 6.47 4.88
QAT 9.25 6.90 8.74 7.13 7.19 5.63 7.02 5.48
GPTQ 8.28 6.74 8.20 6.62 7.24 5.63 7.10 5.56
OminiQuant 7.75 6.03 - - 6.98 5.28 - -
AWQ 7.84 6.24 - - 6.94 5.32 - -
SqueezeLLM 7.51 5.96 - - 6.82 5.23 - -
CherryQ 7.39 5.92 7.34 5.86 6.80 5.26 6.76 5.21
Table 4 extends the comparison to 4-bit quantization. CherryQ continues to excel, achieving the highest scores on most individual tasks and the highest average score overall. These results highlight the generalization ability of CherryQ across different quantization bits and model sizes.
Table 2: Perplexity ($\downarrow$) of 4-bit quantization on LLaMA2 models.
Method 7B-4bit-g128 13B-4bit-g128
c4 wiki2 c4 wiki2
FP16 6.47 4.88 6.47 4.88
QAT 7.29 5.81 6.67 5.12
GPTQ 7.30 5.73 6.63 4.97
OminiQuant 7.12 5.58 6.56 4.95
AWQ 7.19 5.68 6.62 5.05
CherryQ 7.07 5.58 6.56 4.99
Table 3: Performance of different 3-bit quantization methods on Huggingface OpenLLM for LLaMA2-7B and LLaMA2-13B.
Method Hellaswag Winogrande ARC TruthfulQA GSM8K MMLU Average ($\uparrow$)
LLaMA2-7B-3bit-g64
FP16 78.6 74.0 53.2 38.8 14.5 46.7 51.0
QAT 75.5 71.6 49.2 37.3 7.3 40.6 46.9
GPTQ 73.9 71.7 48.6 38.8 8.1 39.4 46.8
CherryQ 77.0 71.8 50.6 38.6 10.4 43.9 48.7
LLaMA2-7B-3bit-g128
FP16 78.6 74.0 53.2 38.8 14.5 46.7 51.0
QAT 75.4 70.8 48.2 37.7 6.7 39.0 46.3
GPTQ 72.9 70.8 48.6 39.1 5.4 38.2 45.8
CherryQ 76.3 72.4 49.7 38.1 8.8 41.6 47.8
LLaMA2-13B-3bit-g64
FP16 82.1 76.6 59.4 37.4 22.5 55.7 55.6
QAT 80.7 75.1 55.5 39.0 16.8 52.9 53.3
GPTQ 79.2 74.4 56.5 36.0 16.4 52.4 52.5
CherryQ 81.1 76.2 57.3 38.0 18.4 53.5 54.1
LLaMA2-13B-3bit-g128
FP16 82.1 76.6 59.4 37.4 22.5 55.7 55.6
QAT 80.7 75.5 55.3 38.8 16.0 51.9 53.0
GPTQ 79.1 75.4 54.1 34.9 15.6 50.3 51.6
CherryQ 81.0 75.4 56.7 38.9 17.8 52.5 53.7
Table 4: Performance comparison of different 4-bit quantization methods for LLaMA2-7B and LLaMA2-13B models over Huggingface OpenLLM Leaderboard.
Method Hellaswag Winogrande ARC TruthfulQA GSM8K MMLU Average ($\uparrow$)
LLaMA2-7B-4bit-g128
FP16 78.6 74.0 53.2 38.8 14.5 46.7 51.0
QAT 77.5 72.2 52.0 39.0 10.6 43.7 49.2
GPTQ 77.6 72.9 52.0 39.1 11.1 43.8 49.4
CherryQ 77.8 73.5 51.5 39.5 12.9 44.4 49.9
LLaMA2-13B-4bit-g128
FP16 82.1 76.6 59.4 37.4 22.5 55.7 55.6
QAT 81.9 75.7 57.9 38.9 19.6 54.2 54.7
GPTQ 81.5 76.8 57.4 36.1 20.4 54.6 54.5
CherryQ 82.0 77.0 58.6 38.8 21.0 54.6 55.3
6.3 Effect of Chat LLM Quantization
We conduct experiments on Vicuna-1.5 [5]. We apply 3-bit quantization with group size=128 for CherryQ and other baselines.
Evaluation To assess the performance of quantized open-ended chat models, we employ a pairwise comparison on the Vicuna-bench [26], which consists of 80 test samples. We compare the responses generated by the quantized models against those generated by the original 16-bit Vicuna-1.5. The evaluation is performed using GPT-4, which automatically classifies the quantized model’s response as “win”, “tie”, or “lose” relative to the FP16 model’s response. To get rid of the ordering effect of the evaluation, we follow [17] to compare the responses with both orders, leading to 160 trials.
Figure 3 presents the results of the pairwise comparison for each quantized model against its FP16 counterpart. The results demonstrate that CherryQ consistently outperforms other quantization baselines in preserving the performance of chat models. It achieves the highest number of wins and ties against the FP16 models, while minimizing the number of losses.
Notably, 3-bit CherryQ achieves a slightly better win-tie-lose ratio over the FP16 Vicuna model, indicating that the 3-bit quantized model performs on par with or even better than the FP16 model. As intuitively CherryQ cannot surpass the target 16 bit model, we think the result suggests that CherryQ maintains almost all its performance even at 3 bit, making GPT-4 hard to distinguish the quality of low-bit and FP16 models.
Figure 3: Comparison of 3-bit quantized models to FP16 Vicuna-1.5. (Left) Comparisons to Vicuna-1.5-7B. (Right) Comparisons to Vicuna-1.5-13B. CherryQ even shows competitive quality compared to the 16-bit counterpart.
6.4 Comparison of Parameter Selection Criteria
To evaluate the effectiveness of our proposed impact-based parameter selection criterion, we conduct experiments comparing it with the commonly used magnitude-based criterion [18]. Table 5 presents the perplexity of LLaMA2-7B-3bit and LLaMA2-13B-3bit models, using both criteria for cherry parameter selection.
From the results, it is evident that the impact-based criterion consistently outperforms the magnitude-based criterion across all settings. These results demonstrate that our proposed impact-based criterion is a more effective measure of parameter importance compared to the magnitude-based criterion. The impacts identify and preserve the most critical parameters during the quantization process. We think this justify the heterogeneity of parameter impacts against parameter magnitudes as we highlighted in § 5.
Table 5: Perplexity ($\downarrow$) of different parameter selection criteria.
Method LLaMA2-7B-3bit LLaMA2-13B-3bit
c4 wiki2 c4 wiki2
Magnitude-g64 7.93 6.40 6.91 5.35
Impact-g64 7.34 5.86 6.76 5.21
Magnitude-g128 8.12 6.58 6.94 5.37
Impact-g128 7.39 5.92 6.80 5.26
LLaMA2-7B-4bit LLaMA2-13B-4bit
Magnitude-g128 7.19 5.68 6.62 5.05
Impact-g128 7.07 5.58 6.56 4.99
The extensive experimental results presented in this section clearly demonstrate the superiority of CherryQ compared to existing quantization methods. By effectively identifying the critical cherry parameters and unifying the mixed-precision parameter optimization, CherryQ achieves state-of-the-art performance for both base LLMs and chat LLMs.
7 Conclusion
In this paper, we investigated the parameter heterogeneity phenomenon in LLMs. Our experiments on LLaMA2, Mistral, Gemma, and Vicuna models, consistently demonstrated that a small subset of parameters plays a crucial role in maintaining the model’s performance, while the majority of parameters can be quantized to ultra-low precision without significant degradation. This finding highlights the potential for efficient model compression and quantization techniques that take into account the heterogeneous nature of parameter importance.
Motivated by this observation, we proposed a novel impact-based parameter selection criterion for quantization. Our method effectively identifies and preserves the most critical cherry parameters during the quantization process. We use a QAT framework for unified optimization of both cherry parameters and normal parameters. Extensive experiments demonstrate that CherryQ outperforms the commonly used magnitude-based criterion, achieving significantly lower perplexity scores and better downstream performance. The heterogeneity and proposed approach pave the way for more efficient deployment of LLMs in resource-constrained environments.
Appendix A Effect of Chat LLM Quantization on MMLU
We further evaluate the performance of CherryQ on the MMLU benchmark by quantizing the Vicuna1.5 model. As shown in Table 6, CherryQ outperforms both QAT and GPTQ in terms of average accuracy across almost all categories.
Table 6: Comparison of different 3-bit quantization methods on zero-shot MMLU accuracy applied to Vicuna-1.5.
Method Humanities STEM Social Sciences Other Average
Vicuna1.5-7B-3bit-g128
FP16 46.8 39.4 57.9 56.3 49.9
QAT 43.4 37.7 53.0 52.4 46.4
GPTQ 42.7 37.3 53.0 51.0 45.7
CherryQ 43.8 37.2 54.3 53.5 46.9
Vicuna1.5-13B-3bit-g128
FP16 50.2 43.5 63.0 62.0 54.3
QAT 47.8 40.1 58.6 58.1 50.9
GPTQ 46.1 39.4 57.6 55.2 49.3
CherryQ 49.0 40.6 60.2 58.8 51.9